# citations.bib

### v1

@@ 5.5 Comparison with Floorplan Generation Approach
As illustrated in figure 12, 360-DFPE generates a 2D layout only for the enclosed scenes, failing to account for partial rooms. Conversely, our approach, leveraging primitive extraction, consistently addresses such scenarios. This distinction is further underscored by the RMSE values presented in table 5. Despite minor imperfections observed in our approach due to significant amount of noise in input point clouds, the overall scene layout is preserved during the wall mesh clipping.
Since our approach relies on the quality of input point clouds for primitives extraction, in scenes with significant noise like EDJbRE in figure 12, our approach may exhibit unwanted wall partitions, as highlighted by the blue region. In contrast, 360-DFPE estimates the floorplan from the wall-ceiling-floor segmentation leveraging HorizonNet [sun2019horizonnet], thus the unwanted wall mesh partitions highlighted by the red region in figure 12, are due to the imperfections during the segmentation phase and not due to the noises in the input point clouds. Comparing the RMSE values in table 5, our approach shows significantly lower RMSE.
@@
Figure 12: Comparison with floorplan estimation approach. The wall mesh, highlighted by the blue region, is generated due to a large noise in the input point cloud and the mesh, highlighted by the red region, is due to the imperfections (such as occlusions, low quality images, poor segmentation by HorizonNet, etc.) in wall-ceiling-floor segmentation.
@@ 8 Acknowledgements
Vertex translation and mesh clipping algorithms–S.R. and M.A.; Axis alignment algorithm and segmentation of structured and non-structured scenes–B.K.; Comparisons against surface reconstruction, shape approximation and floorplan generation approaches–S.R. and B.K.; Validation, Formal analysis, Software, Investigation M.A., B.K. and S.R.; Writing original draft preparation–S.R., B.K., and M.A.; Draft review and editing–S.R., B.K., M. A, and V.O.; Resources, V.O.; Supervision, M.A and V.O; Project administration, M.A and V.O;

### v2

@@ 5.5 Comparison with Floorplan Generation Approach
As illustrated in figure 12, 360-DFPE generates a 2D layout only for the enclosed scenes, failing to account for partial rooms. Conversely, our approach, leveraging primitive extraction, consistently addresses such scenarios. This distinction is further underscored by the RMSE values presented in table 5. Despite minor imperfections observed in our approach due to a significant amount of noise in input point clouds, the overall scene layout is preserved during the wall mesh clipping.
Since our approach relies on the quality of input point clouds for primitive extraction, in scenes with significant noise like EDJbRE in figure 12, our approach may exhibit unwanted wall partitions, as highlighted by the blue region. In contrast, 360-DFPE estimates the floorplan from the wall-ceiling-floor segmentation leveraging HorizonNet [sun2019horizonnet], thus the unwanted wall mesh partitions highlighted by the red region in figure 12, is due to the imperfections during the segmentation phase and not due to the noises in the input point clouds. Comparing the RMSE values in table 5, our approach shows significantly lower RMSE.
@@
Figure 12: Comparison with floorplan estimation approach. The wall mesh, highlighted by the blue region, is generated due to a large noise in the input point cloud and the mesh, highlighted by the red region, is due to the imperfections (such as occlusions, low-quality images, poor segmentation by HorizonNet, etc.) in wall-ceiling-floor segmentation.
@@ 8 Acknowledgements
Vertex translation and mesh clipping algorithms–S.R. and M.A.; Axis alignment algorithm and segmentation of structured and non-structured scenes–B.K.; Comparisons against surface reconstruction, shape approximation, and floorplan generation approaches–S.R. and B.K.; Validation, Formal analysis, Software, Investigation M.A., B.K. and S.R.; Writing original draft preparation–S.R., B.K., and M.A.; Draft review and editing–S.R., B.K., M. A, and V.O.; Resources, V.O.; Supervision, M.A and V.O; Project administration, M.A, and V.O;